OpenStudy (anonymous):

Can some one explain to me how to do this? I don't really understand... Will give medal!!! write an equation in point slope form of the line that passes through point j(4, 1) with slope -4

OpenStudy (anonymous):

For this, you're referring to something called point-slope form. If you have the slope of a line as well as a point on that line, you can find out what that line is by using: \[y-y_{1} = m(x-x_{1})\] In this equation, m is the slope and x1 and y1 are the coordinates of the point we know. Since the problem tells us the slope is -4, then we have m = -4. We are told that our line contains the point (4,1), which means x1 = 4 and y1 = 1. Now all you need to do is plug those 3 things into that point-slope formula I posted.

OpenStudy (anonymous):

So, the answer would be y - 1 = -4(x - 4) ?